OpenStudy (anonymous):

Find the radius and center of the circle. equation: x^2+y^2-8y-4=-16. & there is no just x term. I don't know why..

OpenStudy (lgbasallote):

that means h is 0...

OpenStudy (lgbasallote):

like center is at (0,1) or something..

OpenStudy (anonymous):

first you need to convert it into standard form standard form with center (a,b) and radius r is \[\Large (x-a)^2+(y-b)^2=r^2\] x^2+y^2-8y-4=-16 add and subtract 16 on left side \[\Large x^2+y^2-8y+16-16-4=-16\] simplify \[\Large (x)^2+(y-4)^2=4\] can you tell me now the center and radius by comparing with \[\Large (x-a)^2+(y-b)^2=r^2\]

OpenStudy (anonymous):

Why do you add & subtract 16 on the left side? Isn't that redundant?

OpenStudy (lgbasallote):

he likes to confuse himself ;D

OpenStudy (anonymous):

to make complete square of (y-4)^2 if you expand it you will get the original equation back.

OpenStudy (anonymous):

ok if you are not comfortable with that. can i write x^2+y^2-8y-4=-16. as x^2+y^2-8y+16=4 ??

OpenStudy (anonymous):

The center is (0,4) and the radius is 2. x^2+(y-4)^2=4

OpenStudy (anonymous):

@xThatOneKidx do not provide just answer. you need this. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@sami-21 Okay, well I kind of understand. A little hazy, though.

OpenStudy (anonymous):

Lulz my bad